You will purchase some CDs and DVDs. If you purchase 11 CDs and 9 DVDs, it will cost you $85.35; if you purchase 9 CDs and 11 DVDs, it will cost you $95.65. Solve for how much a CD and a DVD cost.
Answer by lwsshak3(11628)

You will purchase some CDs and DVDs. If you purchase 11 CDs and 9 DVDs, it will cost you $85.35; if you purchase 9 CDs and 11 DVDs, it will cost you $95.65. Solve for how much a CD and a DVD cost.
***
let x=cost of each CD
let y=cost of each DVD
..
11x+9y=85.35
9x+11y=95.65
..
99x+81y=768.15(mult. eq1 by 9)
99x+121y=1052.15(mult. eq2 by 11)
subtract to eliminate x
-40y=-284
y=7.10
..
9x=95.65-11y=17.55
x=1.95
cost of each CD=$1.95
cost of each DVD=$7.10